OpenStudy (anonymous):

Is this correct? Simplify 4k^3(3k^3 – 5k^2 – k + 1) answer... 12k^6 - 20k^5 - 4k^4 + 4k^3 ?

OpenStudy (akshay_budhkar):

Yes it is!

OpenStudy (anonymous):

no it's not

OpenStudy (anonymous):

the question asks you to simplify it, not expand it

OpenStudy (anonymous):

here's a hint: (k-1)(k-1)(3k+1) = (k^2 -2k +1)(3k+1) = 3k^3 -5k^2 -k +1

OpenStudy (anonymous):

so your answer is actually: 4k^3(3k^3 – 5k^2 – k + 1) = 4k^3(k-1)^2(3k+1)

OpenStudy (anonymous):

thats not even one of my options o.O

OpenStudy (anonymous):

well that's what it simplifies to. The thing in brackets is a cubic which you can factor. Simplifying is expressing something in its most compact form...

OpenStudy (anonymous):

hm okay thanks :)